OpenStudy (anonymous):

graph and show work for y+x=-6

OpenStudy (anonymous):

first isolate y from x and numbers

OpenStudy (anonymous):

you will get y=-6-x. now to graph this, first graph y=-x, and then shift it on 6 units down

OpenStudy (anonymous):

y=x+-6

OpenStudy (anonymous):

to isolate do the following: 1) y+x=-6 2) y+x-x=-6-x 3)y=-6-x or y=-x-6

OpenStudy (anonymous):

does it make sense? when u add and subtract equal expressions from both sides the equality stays the same

OpenStudy (anonymous):



OpenStudy (anonymous):

so does that mean that x=-6 also?

OpenStudy (anonymous):